OpenStudy (anonymous):

Slope: When the slope is -2/3, the author says "if we go over by 1, we gonna go down by 2/3rds". How is this possible? What's the mathematical process through which he determines this?

OpenStudy (amistre64):

slope is defined to be the change in y values divided by the change in x values

OpenStudy (anonymous):

I know what is slope, and I know that in -2/3rds when you go over by 2, you go down by 3. But how do you then translate this for smaller values? How do you get from -2/3 to -1 over 2/3?

OpenStudy (amistre64):

-2/3 ----- = (-2/3)/1 3

OpenStudy (amistre64):

instead of moving by -2 and 3; move by 1/3 less

OpenStudy (amistre64):

or, we can simply define it as the equation: y = -2/3 x when x=1, y=-2/3

OpenStudy (anonymous):

isn't -2/3 divided by 3 -2/9? in reply to -2/3 ----- = (-2/3)/1 3

OpenStudy (amistre64):

depends on how you read my notation :/

OpenStudy (amistre64):

divide both top and bottom by 3

OpenStudy (anonymous):

or, we can simply define it as the equation: y = -2/3 x when x=1, y=-2/3 ---------- this made a lot of sense

OpenStudy (amistre64):

i knew something has to :)

OpenStudy (anonymous):

Unfortunately it doesn't seen to make sense yet, in another example when the slope m=3, he says that if x is equal to 0, y is equal to 1. so if we use the previous formula y = 3 * 0 , it means y is zero as well, which doesn't hold true.

OpenStudy (amistre64):

that just means that at x=0, y=1 which is the point at which the line intercepts or cross the y axis

OpenStudy (amistre64):

y = slope*x + yintercept

OpenStudy (amistre64):

there are an infinity of lines that can have the same slope; they are defined specifically by where they cross, intercept, the y axis

OpenStudy (amistre64):

so in this case: y = 3x + c would ba a family of lines; and we want the one that when x=0, y=1 y = 3x + 1 satisfies that requirement

OpenStudy (amistre64):

it is commonly seen as: y = mx + b

OpenStudy (anonymous):

I'm sorry but I didn't get you, can someone else please reply, will try to open a new question.

OpenStudy (amistre64):

the y = 3x still applies; but we are given new information such that when x=0, y=1 y = 3*0 = 0 so we have to adjust this to match the given information; when does 0 equal 1?

OpenStudy (amistre64):



OpenStudy (amistre64):



OpenStudy (anonymous):

http://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/v/graphs-using-slope-intercept-form I'm talking about this specific video, he's talking about finding the m if you have the slope only

OpenStudy (anonymous):

*the b

OpenStudy (amistre64):

you cant find b if you only have a slope; you need to know more about the situation; you need to know at least one other point that the line cross thru, or anchors to

OpenStudy (amistre64):

the b is defined as the point where the line crosses the y axis; its refered to as the y-intercept

OpenStudy (anonymous):

yeah we also know the x,y but he's trying to find the b non-algebrically by guessing somehow